# Aesthetics is Cheap, Show me the Text: An Empirical Evaluation of State-of-the-Art Generative Models for OCR

### v2

@@ Abstract
Text image is a unique and crucial information medium that integrates visual aesthetics and linguistic semantics in modern e-society. Due to their subtlety and complexity, the generation of text images represents a challenging and evolving frontier in the image generation field. The recent surge of specialized image generators (e.g., Flux-series) and unified generative models (e.g., GPT-4o), which demonstrate exceptional fidelity, raises a natural question: can they master the intricacies of text image generation and editing? Motivated by this, we assess current state-of-the-art generative models’ capabilities in terms of text image generation and editing. We incorporate various typical optical character recognition (OCR) tasks into our evaluation and broaden the concept of text-based generation tasks into OCR generative tasks. We select 33 representative tasks and categorize them into five categories: document, handwritten text, scene text, artistic text, and complex & layout-rich text. For comprehensive evaluation, we examine six models across both closed-source and open-source domains, using tailored, high-quality image inputs and prompts. Through this evaluation, we draw crucial observations and identify the weaknesses of current generative models for OCR tasks. We argue that photorealistic text image generation and editing should be internalized as foundational skills into general-domain generative models, rather than being delegated to specialized solutions, and we hope this empirical analysis can provide valuable insights for the community to achieve this goal. This evaluation is online and will be continuously updated at our GitHub repository.
[…]
1 Introduction
@@
Generating images with machines represents humanity’s ambitious pursuit to translate the visual world into algorithms, teaching computers the fundamental human skill of creation. Since the 1960s, researchers have started to use computers for creating artworks like films and still ASCII art [1, 2, 3, 4]. Later, this field evolved from early human-programmed, procedural rules [5, 6] to modern, data-driven artificial intelligence (AI) approaches, exemplified by the early Variational AutoEncoders (VAE) [7, 8] and Generative Adversarial Networks (GAN) [9, 10]. Beyond image-to-image translation or conditional image generation [11], researchers sought to control the generation process more naturally and flexibly. This fueled the emergence of text-to-image (T2I) generation [12], allowing users, even those non-technical, to generate visual content with natural language. Building upon the proliferation of T2I generation [13, 14, 15], the community has now broaden the application realm into instruction-guided image generation and editing, driven by the confluence of large-scale text-image datasets, advances of language models [16, 17], and innovative architectures like diffusion [18, 19], flow matching [20], and autoregressive generation [21, 22].
@@
Mainstream state-of-the-art (SOTA) generation approaches fall into two types: specialized image generation models and native unified generative models. The first type is primarily built based on diffusion or flow matching framework, such as DALL·E [23, 24, 25], Imagen [26], Stable Diffsuion [19, 27, 28], Flux [29, 30], and OminGen [31, 32]. They focus merely on generating or editing images based on instructions, using a relatively small text encoder [16] for textual prompt encoding. In contrast, the second type unifies image understanding and generation within one Transformer model, termed unified generative models. They are typically built based on multimodal large language models (MLLMs) with the autoregressive objective, which can be further divided into three meta-categories. (1) Models w/o diffusion modeling. This line of models use VQGAN [33] or VQVAE [34] to tokenize images into discrete tokens as input to the language models [35, 36] for multimodal autoregressive learning, and detokenizes the output features back into images [37, 38, 39, 40, 41, 42, 43, 44]. (2) Diffusion-embedded models. In contrast, this type of works embrace diffusion modeling, integrating diffusion’s loss and the iterative denosing process into the Transformer for training and inference, respectively [45, 46, 47, 48, 49, 50]. They can switch their role as language models for autoregressive text generation or a diffusion model for parallel image synthesis. (3) Models with extra diffusion head. Compared to the second meta-category, this fashion of approach retains only autoregressive next-token prediction loss for training, while setting up an additional, lightweight diffusion head [22] alongside the MLLM, which is conditioned on the MLLM’s hidden states for image generation [51, 52, 53, 54, 55]. Recently, the release of Gemini-2.0-Flash [56] and GPT-4o [57] marks as a milestone of unified native image generation, exhibiting commercial-level image generation and editing capabilities with non-deteriorated understanding skills.
@@
Figure 1: Task categorization and subordination. We classify 33 OCR generative tasks into five primary categories based on text characteristics: document, handwritten text, scene text, artistic text, and complex & layout-rich (CLR) text. Each primary category encompasses multiple sub-tasks.
@@
Text images, such as documents, scene texts, and handwriting, represent special and critical visual information carriers for human communication. Unlike natural objects or paintings, text images contain structured linguistic information that requires both visual perception and language understanding for proper interpretation. This dual requirement makes the generation and manipulation of text images a particularly challenging frontier in the generative AI community [58, 59]. Despite significant advances in large-scale T2I synthesis, generating high-quality text images remains problematic due to fundamental challenges, such as ensuring precise character formation and legibility across scales, maintaining semantic consistency between text and visual content, and handling diverse layouts and orientations. General-domain generators excel at natural image generation but usually falter in text image generation and editing, and very few of them [28, 57, 60] have formally acknowledged or addressed this critical deficiency. We argue that photorealistic text generation is an important and indispensable ability for achieving artificial general intelligence (AGI). Although numerous models tailored for text image synthesis have been proposed [61, 62, 63, 64, 65, 66], text generation should be internalized as a foundational skill, rather than delegated to specialized solutions. In addition, a comprehensive assessment dedicated to text image generation/editing is lacking. Such assessment is essential for systematically evaluating the strengths and weaknesses of existing models, illuminating clear pathways for future research.
@@
To this end, we present a comprehensive evaluation of text image generation and editing capabilities in SOTA generative models. In literature, the automatic interpretation of text images is known as optical character recognition (OCR) [67], among which many tasks are inherently generative tasks, such as document deshadowing [68], scene text removal [69], and font generation [65]. Hence, we broaden the definition of text generation tasks and reframe them as OCR generative tasks. Our evaluation operates in three stages. First, we curate 33 OCR tasks and classify them into five main categories based on text properties: document, handwritten text, scene text, artistic text, and complex & layout-rich (CLR) text. Second, we selected a representative suite of models for benchmarking, including both specialized generators and unified models from closed-source and open-source domains: GPT-4o [57], Qwen-VLo-Preview [60], Flux.1-Kontext-dev [30], OmniGen2 [31], BAGEL [50], and Janus-4o [44]. Third, we tailor expressive prompts and select high-quality input images from public datasets [70, 71] or web sources for each task, encompassing bilingual evaluation in both English and Chinese. From the evaluation, we discover that text image generation and editing remain substantial challenges for existing generative models, and draw several crucial observations as follows.
  1. 1.
@@
  9. 9.
    To-be-improved vertical-domain task performance. For very specialized vertical-domain tasks, such as historical document restoration, historical image style transfer, and modern document tasks, these general-domain models demonstrate inferior proficiency. Incorporating these tasks into model training could improve both general and specialized generation capabilities.
  10. 10.
    Limited multilingual proficiency. While current models are capable of English text generation, some of them, particularly open-source ones, fail to generate Chinese text. The resulting text is often garbled and illegible, with frequent character omissions. Additionally, Chinese text specified in instructions is sometimes incorrectly converted into other languages. Adding multilingual support is essential for universally and linguistically applicable generative models.
  11. 11.
    Large gap between open-source and closed-source models. In terms of text generation and editing, closed-source models, especially GPT-4o, significantly outperform the open-source ones. While this could be limited by the smaller size of open-source models (only 7B), more future efforts should be devoted to improving their text synthesis skills. Scaling laws specifically for text image generation have not been unveiled, representing a promising research direction.
While some prior works [72, 73] have evaluated GPT-4o’s native generation abilities in general object domains, we transition the focus to text images and expand our evaluation to both open-source and closed-source unified generative models. We hope this work provides valuable insights into the strengths and limitations of current generative models, benefiting the design of future models toward better text synthesis abilities.
2 Evaluation Setting
@@
Tested models. Our evaluation considers both closed-source models and open-source models. Closed-source models include GPT-4o [57] and Qwen-VLo-Preview [60]; while open-source models include Flux.1-Kontext-dev [30], OmniGen2 [31], BAGEL [50], and Janus-4o [44]. Except for Flux.1-Kontext-dev and OmniGen2, which are specialized image generation models, all others are unified generative models that are capable of both understanding and generation. In addition, all models support text-to-image (T2I) generation and instruction-based image editing, competent for our evaluation tasks. An overview of these models is summarized in Table 1.
@@
Table 1: Overview of tested models. Pub. denotes publication. Und. denotes understanding. Gen. denotes Generation. GPU Memory specifically refers to GPU memory requirements for model inference.
@@
GPT-4o [57] 2025.3.25 Unified Und. & Gen. Closed-Source - - -
Qwen-VLo-Preview [60] 2025.6.26 Unified Und. & Gen. Closed-Source - - -
Flux.1-Kontext-dev [30] 2025.6.17 Specialized Gen. Open-Source 12B 77 tokens 48GB
OmniGen2 [31] 2025.6.23 Specialized Gen. Open-Source 7B 128,000 tokens 48GB
BAGEL [50] 2025.5.20 Unified Und. & Gen. Open-Source 7B 32,768 tokens 60GB
Janus-4o [44] 2025.6.22 Unified Und. & Gen. Open-Source 7B 4,096 tokens 48GB
3 Result of Document Image
3.1 Modern Document
@@
Figure 2: Results of the document dewarping task. The red glowing indicates the relatively best output corresponding to each input. The overall dewarping results are inferior, where GPT-4o rectifies the image to be flat but loses some embedded graphics and text, while other methods even fail to perform dewarping and lose substantial textual details.
@@
Modern documents plays an essential role in human daily activities for information transmission, consisting of text, graphics, tables, and other informational elements. Therefore, document digitization and processing has become increasingly crucial for efficient information management and automated analysis, with OCR as a core technique [74, 75]. Typically, before performing OCR, the digitized documents should undergo preprocessing to enhance OCR accuracy, such as flattening the scrumpling of the documents, removing shadows, and sharpening the text. These preprocessing technologies constitute different document processing tasks, and most of them can be formulated in an image-to-image translation manner. Given the recent trend of instruction-based image generation and editing, we attempt to investigate whether these models can address important and practical document-related tasks.
@@
Figure 3: Results of the document deshadowing and document deblurring tasks. Flux.1-Kontext-dev exhibits the optimal deshadowing results in both language cases, with most texts preserved and shadow removed. Other models either fail to remove the shadow or mistakenly repeat the textual content. For document deblurring, Flux.1-Kontext-dev showcase nearly perfect result in the English scenario. GPT-4o performs better in the Chinese case with precise text restoration, but fails to restore the document structure.
3.1.1 Document Dewarping
Document dewarping refers to correcting geometric distortions of document images, typically caused by curved, warped, or folded pages when scanned or photographed. Models are required to output a flat document surface with original text preserved, facilitating reading or further OCR processing [76, 77, 78, 79, 80]. Existing dewarping methods compute a displacement field that maps the input, warped image to the output, dewarped image. Yet, from the perspective of image generation, it can be formulated as an image-to-image translation task. Early generative models can not interpret the “dewarping” or “rectification” instructions and handle the drastic and detailed pixel changes to derive an accurate, flat document. Given the advancement of image generators, especially the unified generative models in both understanding and generation, they may now be able to understand the “dewarping” instruction and conduct sophisticated pixel transformation. More broadly, we think this task somehow represents the perception and comprehension abilities of models to the real physics world and therefore conduct a detailed evaluation on this task.
Results of the document dewarping task from different models are presented in Fig. 2. As observed, all the dewarped results are unsatisfactory. GPT-4o is the only model that identifies the instruction of “dewarping” and delivers flattened output documents, showcasing impressive instruction-following capability. While this is a surprising discovery, GPT-4o loses a large portion of original text content and some embedded graphics, also fails to render Chinese content, making the dewarped output infeasible for real-world usage. For other models, they mostly fall short in understanding and following the “dewarping” or “crop” instruction, and generate blurred, chaotic text. Qwen-VLo-Preview even exhibits hallucination without following the input and instructions, as shown in the penultimate row, third column of Fig. 2.
@@ 3.1.1 Document Dewarping
Figure 4: Results of the document appearance enhancement task. Only GPT-4o and Flux.1-Kontext-dev are capable of comprehending the instruction of outputting “PDF-like” documents. However, most of them fall short in preserving document structures, particularly demonstrated by the mistaken repetition of table contents.
Figure 5: Text editing results for modern document. The first three models can follow the instruction (at least partially) while the last three models can not, potentially due to their smaller parameter scale (only 7B). This suggests that document text editing can be addressed by existing generative models when sufficient model scale is available.
3.1.2 Document Deshadowing and Document Deblurring

### v3

@@ Abstract.
Text image is a unique and crucial information medium that integrates visual aesthetics and linguistic semantics in modern e-society. Due to their subtlety and complexity, the generation of text images represents a challenging and evolving frontier in the image generation field. The recent surge of specialized image generators (e.g., Flux-series) and unified generative models (e.g., GPT-4o), which demonstrate exceptional fidelity, raises a natural question: can they master the intricacies of text image generation and editing? Motivated by this, we assess current state-of-the-art generative models’ capabilities in terms of text image generation and editing. We incorporate various typical optical character recognition (OCR) tasks into our evaluation and broaden the concept of text-based generation tasks into OCR generative tasks. We select 33 representative tasks and categorize them into five categories: document, handwritten text, scene text, artistic text, and complex & layout-rich text. For comprehensive evaluation, we examine seven models across both closed-source and open-source domains, using tailored, high-quality image inputs and prompts. Through this evaluation, we draw crucial observations and identify the weaknesses of current generative models for OCR tasks. We argue that photorealistic text image generation and editing should be internalized as foundational skills into general-domain generative models, rather than being delegated to specialized solutions, and we hope this empirical analysis can provide valuable insights for the community to achieve this goal. This evaluation is online at https://github.com/NiceRingNode/Awesome-Generative-Models-for-OCR.
OCR, Generative Models, Unified Understanding and Generation
$c$s: Human-centered computing Empirical studies in visualization
1. Introduction
@@
images with machines represents humanity’s ambitious pursuit to translate the visual world into algorithms, teaching computers the fundamental human skill of creation. Since the 1960s, researchers have started to use computers for creating artworks like films and still ASCII art (Noll, 2016; Krampen and Seitz, 1967; Goodchild, 2021; Computer History Museum, 2024). Later, this field evolved from early human-programmed, procedural rules (Cohen, 1973; Perlin, 1985) to modern, data-driven artificial intelligence (AI) approaches, exemplified by the early Variational AutoEncoders (VAE) (Kingma et al., 2013; Higgins et al., 2017) and Generative Adversarial Networks (GAN) (Goodfellow et al., 2014; Karras et al., 2019). Beyond image-to-image translation or conditional image generation (Elasri et al., 2022), researchers sought to control the generation process more naturally and flexibly. This fueled the emergence of text-to-image (T2I) generation (Reed et al., 2016), allowing users, even those non-technical, to generate visual content with natural language. Building upon the proliferation of T2I generation (Qiao et al., 2019; Li et al., 2019; Nonghai Zhang, 2024), the community has now broaden the application realm into instruction-guided image generation and editing, driven by the confluence of large-scale text-image datasets, advances of language models (Radford et al., 2021; Raffel et al., 2020), and innovative architectures like diffusion (Ho et al., 2020; Rombach et al., 2022), flow matching (Lipman et al., 2023), and autoregressive generation (Chen et al., 2020; Li et al., 2024a).
@@
Mainstream state-of-the-art (SOTA) generation approaches fall into two types: specialized image generation models and native unified generative models. The first type is primarily built based on diffusion or flow matching framework, such as DALL·E (Ramesh et al., 2021, 2022; OpenAI, 2023), Imagen (Saharia et al., 2022), Stable Diffsuion (Rombach et al., 2022; Podell et al., 2023; Esser et al., 2024), Flux (Labs, 2025; Labs et al., 2025), and OminGen (Wu et al., 2025d; Xiao et al., 2025). They focus merely on generating or editing images based on instructions, using a relatively small text encoder (Radford et al., 2021) for textual prompt encoding. In contrast, the second type unifies image understanding and generation within one Transformer model, termed unified generative models. They are typically built based on multimodal large language models (MLLMs) with the autoregressive objective, which can be further divided into three meta-categories. (1) Models w/o diffusion modeling. This line of models uses VQGAN (Esser et al., 2021) or VQVAE (van den Oord et al., 2017) to tokenize images into discrete tokens as input to the language models for multimodal autoregressive learning, and detokenizes the output features back into images (Lu et al., 2024; Team, 2024; Liu et al., 2025; Wu et al., 2025a; Ma et al., 2025; Wu et al., 2024; Wang et al., 2024; Chen et al., 2025b). (2) Diffusion-embedded models. In contrast, this type of works embrace diffusion modeling, integrating diffusion’s loss and the iterative denosing process into the Transformer for training and inference, respectively (Zhou et al., 2024; Xie et al., 2025a, b; Zhao et al., 2024; Shi et al., 2024; Deng et al., 2025). They can switch their role as language models for autoregressive text generation or diffusion models for parallel image synthesis. (3) Models with extra diffusion head. Unlike the second meta-category, this fashion of approach retains only autoregressive next-token prediction loss for training while incorporating an additional, lightweight diffusion head (Li et al., 2024a) alongside the MLLM, which is conditioned on the MLLM’s hidden states for image generation (Fan et al., 2025; Tong et al., 2024; Sun et al., 2024; Pan et al., 2025; Wu et al., 2025c). Recently, the release of Gemini-2.0-Flash (DeepMind, 2025) and GPT-4o (OpenAI, 2025) marks a milestone of unified native image generation, exhibiting commercial-level image generation and editing capabilities without compromising understanding skills.
@@
Figure 1. Task categorization and subordination. We classify 33 OCR generative tasks into five primary categories based on text characteristics: document, handwritten text, scene text, artistic text, and complex & layout-rich (CLR) text. Each primary category encompasses multiple sub-tasks.
@@
Text images, such as documents, scene texts, and handwriting, represent special and critical visual information carriers for human communication. Unlike natural objects or paintings, text images contain structured linguistic information that requires both visual perception and language understanding for proper interpretation. This dual requirement makes the generation and manipulation of text images a particularly challenging frontier in the generative AI community (Elanwar and Betke, 2025; Ren et al., 2025). Despite significant advances in large-scale T2I synthesis, generating high-quality text images remains problematic due to fundamental challenges, such as ensuring precise character formation and legibility across scales, maintaining semantic consistency between text and visual content, and handling diverse layouts and orientations. General-domain generators excel at natural image generation but usually falter in text image generation and editing, and very few of them (Esser et al., 2024; OpenAI, 2025; Team, 2025) have formally acknowledged or addressed this critical deficiency. We argue that photorealistic text generation is an important and indispensable ability for achieving artificial general intelligence (AGI). Although numerous models tailored for text image synthesis have been proposed (Chen et al., 2023; Tuo et al., 2024; Zhu et al., 2023; Luo et al., 2023; Yang et al., 2024; Wang et al., 2025), text generation should be internalized as a foundational skill, rather than delegated to specialized solutions. In addition, despite numerous evaluation on diverse content generation fields, such as face/code/human motion generation, (Kammoun et al., 2022; Schäfer et al., 2024; Zhu et al., 2024a; Fakhoury et al., 2024), a comprehensive assessment dedicated to text image generation/editing is lacking. Such an assessment is essential for systematically evaluating the strengths and weaknesses of existing models, illuminating clear pathways for future research.
@@
To this end, we conduct a comprehensive evaluation of text image generation and editing capabilities in SOTA generative models. In literature, the automatic interpretation of text images is known as optical character recognition (OCR) (Islam et al., 2017), among which many tasks are inherently generative, such as document deshadowing (Zhang et al., 2024b), scene text removal (Liu et al., 2020b), and font generation (Yang et al., 2024). Hence, we broaden the scope of text generation and reframe these problems as OCR generative tasks. Our evaluation operates in three stages. First, we curate 33 OCR tasks and classify them into five main categories based on text properties: document, handwritten text, scene text, artistic text, and complex & layout-rich (CLR) text. Second, we select several representative models for benchmarking, including both specialized generators and unified models from closed-source and open-source domains: GPT-4o (OpenAI, 2025), Qwen-VLo-Preview (Team, 2025), Qwen-Image (Wu et al., 2025b), Flux.1-Kontext-dev (Labs et al., 2025), OmniGen2 (Wu et al., 2025d), BAGEL (Deng et al., 2025), and Janus-4o (Chen et al., 2025b). Third, we tailor expressive prompts and select high-quality input images from public datasets (Chng et al., 2019; Hradiš et al., 2015) or the web for each task, encompassing bilingual evaluation in both English and Chinese. From the evaluation, we discover that text image generation and editing remain substantial challenges for existing generative models, and draw several crucial observations as follows.
  1. (1)
@@
  9. (9)
    To-be-improved vertical-domain task performance. For highly specialized vertical-domain tasks, such as modern document processing tasks and historical document restoration, these general-domain models demonstrate inferior proficiency. Incorporating these tasks into model training could improve both general and specialized generation capabilities.
  10. (10)
    Limited multilingual proficiency. While current models handle English text generation, many fail to generate Chinese text, particularly open-source ones. The generated Chinese text is often garbled, illegible, or missing characters, and instructions specifying Chinese text are sometimes incorrectly rendered in other languages. Adding multilingual support is essential for universally and linguistically applicable generative models.
  11. (11)
    Large gap between open-source and closed-source models. In terms of text generation and editing, closed-source models, especially GPT-4o, significantly outperform the open-source ones. While this could be partially limited by the smaller size of open-source models (7B), more future efforts should be devoted to improving their text synthesis skills. Scaling laws specifically for text image generation have not been explicitly unveiled, representing a promising research direction.
Notably, while some prior works (Chen et al., 2025a; Cao et al., 2025; Li et al., 2025) have evaluated closed-source unified generative models’ (particularly GPT-4o) generation abilities in general visual domains, we transition the focus to text images and expand our evaluation to both open-source and closed-source models. We hope this work can inspire broader research efforts toward improving visual text generation and editing, establishing these skills as fundamental capabilities rather than specialized niches.
2. Evaluation Setting
@@
Tested models. Our evaluation considers both closed-source models and open-source models. Closed-source models include GPT-4o (OpenAI, 2025) and Qwen-VLo-Preview (Team, 2025); while open-source models include Qwen-Image (Wu et al., 2025b), Flux.1-Kontext-dev (Labs et al., 2025), OmniGen2 (Wu et al., 2025d), BAGEL (Deng et al., 2025), and Janus-4o (Chen et al., 2025b). Except for Flux.1-Kontext-dev and OmniGen2, which are specialized image generation models, all others are unified generative models that are capable of both understanding and generation. In addition, all models support text-to-image (T2I) generation and instruction-based image editing, competent for our evaluation tasks. An overview of these models is summarized in Table 1.
@@
Table 1. Overview of tested models. Pub. denotes publication. Und. denotes understanding. Gen. denotes Generation. GPU Memory specifically refers to GPU memory requirements for model inference.
@@
GPT-4o (OpenAI, 2025) 2025.3.25 Unified Und. & Gen. Closed-Source - - -
Qwen-VLo-Preview (Team, 2025) 2025.6.26 Unified Und. & Gen. Closed-Source - - -
Qwen-Image (Wu et al., 2025b) 2025.8.18 Specialized Gen. Open-Source 20B 32,768 tokens 3 $\times$ 48GB
Flux.1-Kontext-dev (Labs et al., 2025) 2025.6.17 Specialized Gen. Open-Source 12B 77 tokens 48GB
OmniGen2 (Wu et al., 2025d) 2025.6.23 Specialized Gen. Open-Source 7B 128,000 tokens 48GB
BAGEL (Deng et al., 2025) 2025.5.20 Unified Und. & Gen. Open-Source 7B 32,768 tokens 2 $\times$ 48GB
Janus-4o (Chen et al., 2025b) 2025.6.22 Unified Und. & Gen. Open-Source 7B 4,096 tokens 48GB
3. Result of Document Image
3.1. Modern Document
@@
Figure 2. Dewarping results for modern documents. The red glowing indicates the relatively best output corresponding to each input. The overall dewarping results are inferior, where GPT-4o rectifies the image to be flat but loses some embedded graphics and text, while other methods even fail to perform dewarping and lose substantial textual details.
@@
Modern documents play an essential role in daily human activities for information transmission, consisting of text, graphics, tables, and other informational elements. Hence, document digitization and processing have become increasingly crucial for efficient information management and automated analysis, with OCR serving as a core technique (Tang et al., 1996; Subramani et al., 2020). Typically, the digitized documents require preprocessing before OCR to enhance accuracy, including flattening the scrumpling of the documents, removing shadows, and sharpening the text. These preprocessing technologies constitute different document processing tasks, most of which can be formulated as an image-to-image translation manner. Given the recent trend of instruction-based image generation and editing, we investigate whether these models can address important and practical document-related tasks.
@@
Figure 3. Deshadowing and deblurring results for modern documents. Flux.1-Kontext-dev exhibits the optimal deshadowing results in both language cases, with most texts preserved and shadow removed. Other models either fail to remove the shadow or mistakenly repeat the textual content. For document deblurring, Flux.1-Kontext-dev showcases nearly perfect results in the English scenario. GPT-4o performs better in the Chinese case with precise text restoration but fails to restore the document structure.
@@
Figure 4. Appearance enhancement results for modern documents. Only GPT-4o and Flux.1-Kontext-dev successfully interpret the instruction of outputting “PDF-like” documents. Most models fail to preserve document structures, particularly evident in erroneous repetition of table contents.
3.1.1. Document Dewarping
Document dewarping refers to correcting geometric distortions of document images, typically caused by curved, warped, or folded pages during scanning or photographing. Models are required to output a flattened document surface with preserved original text, facilitating reading or further OCR processing (Zhang et al., 2022b; Feng et al., 2025; Zhang et al., 2025b, 2026). Existing dewarping methods compute displacement fields that map the warped input image to the dewarped outputs. Yet, from the image generation perspective, this can be formulated as an image-to-image translation task. Early generative models could not interpret the “dewarping” or “rectification” instructions nor handle the drastic, detailed pixel changes to derive an accurate, flattened document. With advances in image generators, especially unified generative models capable of both understanding and generation, they may now understand the “dewarping” instruction and conduct sophisticated pixel transformation. More broadly, this task somehow represents the perception and comprehension abilities of real-world physics, motivating our detailed evaluation.
Results of the document dewarping task are presented in Fig. 2. All the dewarped results prove unsatisfactory. GPT-4o is the only model that identifies the “dewarping” instruction and produces flattened output documents, showcasing impressive instruction-following capability. While this is surprising, GPT-4o loses substantial portions of original text content and some embedded graphics, and also fails to render Chinese content, making the dewarped output infeasible for real-world usage. Other models mostly fail to understand and follow the “dewarping” or “crop” instructions, generating blurred, chaotic text. Qwen-VLo-Preview even exhibits hallucination without following the input and instructions, as shown in the penultimate row, third column of Fig. 2.
Figure 5. Text editing results for modern documents. GPT-4o and Qwen-VLo-preview better follow the instructions compared to other unified models, especially in locating target text and conducting replacement. Flux.1-Kontext-dev significantly outperforms other specialized generation models.
3.1.2. Document Deshadowing
Document deshadowing refers to removing shadows from document images (Li et al., 2023a). Similar to document dewarping, we require models to output deshadowed or deblurred images according to the input image and instructions. Results are presented in the top panel of Fig. 3. For deshadowing, Flux.1-Kontext-dev demonstrates optimal performance by effectively removing shadows while preserving original text content and background color. Other models either fail to remove the shadow (Qwen-Image, OmniGen2, BAGEL), inadvertently repeat text content, or alter the document’s background color (GPT-4o, Qwen-VLo-Preview). Qwen-Image even delivers a mosaic-like output rather than a document as shown in the first row.
3.2. Document Deblurring